Database last updated on Mon, 02 Feb 2026 00:40:50 +0000 by localhost

Instrument constant changes for the week before 2023313 (2 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 15 rows)

Measurements and tests summary.


_jday_loc_serial
313TORONTO#085
(1 sec)

Dead time (0 sec)

Nothing to report.

Standard lamp R6 (1 sec)

Nothing to report.

Standard lamp R5 (2 sec)

Nothing to report.

Number of days since last dome/window cleaning (0 sec, 3 rows)

serialdays
085more than a year
8more than a year

Missing files (0 sec, 1 rows)

Seriallast datadays ago
08520233140-1

Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (1 sec)

No ozone measurements to display.

Unsorted comments (0 sec, 10123 rows)

The END (0 sec, 2 rows)

If you have questions and/or comments about this page please contact Volodya Savastiouk